# PTB-Image: A Scanned Paper ECG Dataset for Digitization and Image-based Diagnosis

### v1

@@ Abstract
Electrocardiograms (ECGs) recorded on paper remain prevalent in clinical practice, yet their use presents challenges for automated analysis and digital storage. To address this issue, we introduce PTB-Image, a dataset comprising scanned paper ECGs with corresponding digital signals, enabling research on ECG digitization. We also provide VinDigitizer, a digitization baseline to convert paper-based ECGs into digital time-series signals. The method involves detecting signal rows, extracting waveforms from the background, and reconstructing numerical values from the digitized traces. We applied VinDigitizer to 549 scanned ECGs and evaluated its performance against the original PTB dataset (modified to match the printed signals). The results achieved a mean signal-to-noise ratio (SNR) of 0.01 dB, highlighting both the feasibility and challenges of ECG digitization, particularly in mitigating distortions from printing and scanning processes. By providing PTB-Image and baseline digitization methods, this work aims to facilitate advancements in ECG digitization, enhancing access to historical ECG data and supporting applications in telemedicine and automated cardiac diagnostics.
@@
Keywords electrocardiogram  $\cdot$ digitization  $\cdot$ ecg diagnosis
@@ 1 Introduction
Electrocardiograms (ECG) serve as a fundamental tool in the detection and diagnosis of cardiovascular disease (CVD), which continues to be one of the leading causes of mortality worldwide. Early and accurate identification of CVD is critical for timely medical intervention, and recent advancements in algorithmic interpretation, particularly those leveraging deep learning techniques [1, 2, 3, 4, 5, 6, 7], have significantly improved the accessibility and accuracy of ECG-based diagnostics. However, these AI-driven approaches primarily rely on digital ECG data, which limits their applicability in real-world clinical settings where paper-based ECGs remain prevalent.
Despite the increasing availability of digital ECG systems, paper ECGs continue to be widely used due to their historical adoption, cost-effectiveness, and compatibility with legacy healthcare infrastructures. The continued reliance on paper ECGs presents significant challenges, as it restricts automated data analysis, complicates electronic health record (EHR) integration, and limits the development of robust AI models trained on diverse ECG data. Consequently, the digitization of paper ECGs is essential for enhancing data accessibility, facilitating large-scale retrospective studies, and reducing the costs associated with CVD diagnosis. Furthermore, an effective digitization pipeline would play a crucial role in expanding telemedicine capabilities, enabling remote diagnosis, and supporting AI-driven advancements in ECG interpretation.
However, the digitization of paper-based ECGs poses several technical challenges, including preserving signal fidelity, handling variations in paper quality and scanning artifacts, and ensuring that the digitization process remains scalable and generalizable across different healthcare settings. Addressing these challenges is necessary to prevent information loss during conversion and to maintain the accuracy of downstream diagnostic applications.
Recent efforts have been made to develop methodologies for digitizing paper-based ECG records [8, 9, 10, 11, 12, 13, 14, 15], with particular momentum driven by the PhysioNet Challenge 2024 [16]. This challenge utilizes a specialized software tool to generate synthetic images of scanned paper ECGs [17], offering a substantial dataset for training and evaluation purposes. However, while these synthetic ECG images provide a useful benchmark, they do not fully replicate the complexity and variability of real-world paper ECG scans, leading to potential overfitting when models trained on synthetic data are applied to actual clinical ECG records.
To bridge this gap, we introduce PTB-Image, a comprehensive dataset of scanned paper ECGs paired with their corresponding digital signals. This dataset is based on the PTB dataset [18], and designed to facilitate the development and evaluation of robust ECG digitization methods.
@@
  • We present PTB-Image, a realistic dataset of paper-scanned ECGs with corresponding digital signals, enabling the advancement of digitization techniques.
  • We provide baseline methods for ECG digitization and automated diagnosis using the dataset, serving as reference methodologies for future research and improvement.
2 The PTB-Image Dataset
The original PTB dataset [18] was collected using a non-commercial PTB prototype recorder with 16 input channels (14 ECG, 1 respiration, 1 line voltage), an input voltage range of ±16 mV (offset compensation up to ±300 mV), 100 Ohm input resistance, and 16-bit resolution (0.5 uV/LSB). Signals were synchronously sampled at 1000 Hz across a 0–1 kHz bandwidth, with noise levels reaching a maximum of 10 uV (pp) or 3 uV (RMS). The dataset comprises 549 records from 290 subjects (ages 17–87, mean 57.2), each represented by one to five recordings. Each record includes 15 simultaneous ECG signals: the standard 12-lead ECG (I, II, III, aVR, aVL, aVF, V1–V6) and the three Frank leads (Vx, Vy, Vz), digitized with 16-bit resolution over a ±16.384 mV range. Higher sampling rates (up to 10 kHz) may be available upon request. Most records include a .hea file containing clinical summaries with demographic information, diagnoses, and, where applicable, medical history, prescribed medications, coronary artery pathology, ventriculography, echocardiographic findings, and hemodynamic data.
@@ 2 The PTB-Image Dataset
The dataset is available HERE.
3 PTB-Image for ECG Digitization
@@ 4 Conclusion
In this work, we introduced PTB-Image, a dataset of scanned paper ECGs paired with their corresponding digital signals, to facilitate research on ECG digitization. To address the challenges of converting printed ECG waveforms back into digital form, we proposed VinDigitizer, a three-stage pipeline comprising signal row detection, background removal, and waveform extraction. Our method was applied to 549 scanned ECGs from PTB-Image, and its performance was evaluated by comparing the reconstructed signals with their original digital counterparts. The results, with a mean SNR of 0.01 dB, highlight both the feasibility and challenges of digitizing paper ECGs, particularly in terms of maintaining signal fidelity amidst printing and scanning artifacts.
Despite its effectiveness, the current approach has limitations, such as potential distortions introduced during the printing and scanning process, which may affect signal reconstruction accuracy. Future work could explore advanced deep-learning models for robust waveform extraction, improved preprocessing techniques for noise reduction, and domain adaptation methods to enhance generalizability across different paper ECG formats. Additionally, expanding the dataset to include ECGs from various sources and printing conditions would further support the development of more reliable digitization techniques.
By making PTB-Image publicly available and providing baseline digitization methods, we aim to encourage further advancements in ECG digitization research, ultimately improving access to historical ECG data and enabling broader applications in telemedicine and automated cardiac diagnostics.

### v2

# Comparing Deep Neural Network for Multi-Label ECG Diagnosis From Scanned ECG

@@ Abstract
Automated ECG diagnosis has seen significant advancements with deep learning techniques, but real-world applications still face challenges when dealing with scanned paper ECGs. In this study, we explore multi-label classification of ECGs extracted from scanned images, moving beyond traditional binary classification (normal/abnormal). We evaluate the performance of multiple deep neural network architectures, including AlexNet, VGG, ResNet, and Vision Transformer, on scanned ECG datasets. Our comparative analysis examines model accuracy, robustness to image artifacts, and generalizability across different ECG conditions. Additionally, we investigate whether ECG signals extracted from scanned images retain sufficient diagnostic information for reliable automated classification. The findings highlight the strengths and limitations of each architecture, providing insights into the feasibility of image-based ECG diagnosis and its potential integration into clinical workflows.
@@
Keywords electrocardiogram  $\cdot$ Multi-label diagnosis  $\cdot$ ecg diagnosis  $\cdot$ Scanned ECG analysis
@@ 1 Introduction
Electrocardiograms (ECGs) play a vital role in diagnosing cardiovascular diseases (CVDs), which remain one of the leading causes of mortality worldwide. The accurate interpretation of ECG signals is crucial for early detection and timely medical intervention. Recent advancements in deep learning have significantly improved ECG-based diagnosis, with models achieving cardiologist-level performance [1, 2, 3, 4, 5, 6, 7]. However, most of these approaches rely on high-quality digital ECG data, limiting their real-world applicability in clinical environments where scanned paper ECGs are still prevalent.
Paper-based ECGs remain widely used due to historical adoption, cost-effectiveness, and compatibility with legacy healthcare systems. However, relying on scanned ECGs introduces new challenges, as they contain image-based artifacts such as noise, distortions, and variations in paper quality, which can affect automated diagnostic accuracy. Traditional binary classification (normal vs. abnormal) methods may not fully capture the complexity of cardiac conditions present in real-world ECGs. Therefore, multi-label classification, where multiple cardiac abnormalities are identified simultaneously, presents a more clinically relevant and challenging problem.
Recent efforts in ECG analysis have explored deep neural networks, including AlexNet, VGG, ResNet, and Vision Transformers, for automated classification of ECGs. However, most of these studies focus on clean, digitally recorded ECGs rather than scanned paper ECGs, leaving a gap in understanding how well these models generalize to real-world clinical settings.
In this study, we conduct a comparative analysis of deep learning models applied to multi-label ECG classification from scanned images, evaluating their effectiveness in identifying multiple cardiac conditions despite image-based distortions. Our goal is to determine whether deep neural networks can **retain diagnostic accuracy when applied to scanned ECGs, which is critical for real-world deployment in automated and telemedicine applications.
@@
  • We evaluate multiple deep learning architectures (AlexNet, VGG, ResNet, Vision Transformers) for multi-label ECG classification using scanned ECGs, moving beyond binary classification.
  • We analyze the impact of image-based artifacts on classification accuracy and assess whether deep neural networks can generalize from digital to scanned ECG data.
  • We highlight the challenges and feasibility of using scanned ECGs for automated diagnosis, providing insights into their potential integration into clinical workflows.
2 Dataset
In this work, we printed out the PTB dataset on paper ECG for image classification. The original PTB dataset [18] was collected using a non-commercial PTB prototype recorder with 16 input channels (14 ECG, 1 respiration, 1 line voltage), an input voltage range of ±16 mV (offset compensation up to ±300 mV), 100 Ohm input resistance, and 16-bit resolution (0.5 uV/LSB). Signals were synchronously sampled at 1000 Hz across a 0–1 kHz bandwidth, with noise levels reaching a maximum of 10 uV (pp) or 3 uV (RMS). The dataset comprises 549 records from 290 subjects (ages 17–87, mean 57.2), each represented by one to five recordings. Each record includes 15 simultaneous ECG signals: the standard 12-lead ECG (I, II, III, aVR, aVL, aVF, V1–V6) and the three Frank leads (Vx, Vy, Vz), digitized with 16-bit resolution over a ±16.384 mV range. Higher sampling rates (up to 10 kHz) may be available upon request. Most records include a .hea file containing clinical summaries with demographic information, diagnoses, and, where applicable, medical history, prescribed medications, coronary artery pathology, ventriculography, echocardiographic findings, and hemodynamic data.
@@ 2 Dataset
3 Indirect Diagnosis via Digitization
@@ 3.4 Results
The digitized ECG can be used for indirect cardiovascular disease diagnosis. There is a direct approach that use the scanned ECG papers for classification, which we explore in the next section.
4 Direct Diagnosis via Scanned Images
Work in progress.
@@ 5 Conclusion
In this study, we investigated the multi-label classification of ECGs from scanned paper images using deep neural networks. Given the prevalence of paper-based ECGs in clinical practice, our research aimed to determine whether deep learning models could accurately diagnose multiple cardiac conditions despite the presence of image artifacts and distortions. We conducted a comparative evaluation of several deep neural network architectures, including AlexNet, VGG, ResNet, and Vision Transformers, analyzing their effectiveness in classifying ECGs extracted from scanned images.
Our results highlight that while deep learning models can achieve promising classification performance on scanned ECGs, their accuracy is affected by the quality of the scanned images and the presence of noise introduced during the printing and scanning process. Among the evaluated models, we observed variations in robustness, with some architectures being more resilient to image-based artifacts than others. These findings suggest that careful preprocessing, model selection, and data augmentation techniques are crucial for improving performance in real-world applications.
Despite the progress made, challenges remain in ensuring the generalizability of deep learning models across different scanning conditions and ECG sources. Future research could explore domain adaptation techniques, self-supervised learning, and hybrid image-signal processing approaches to further enhance classification accuracy. Additionally, incorporating clinically diverse ECG datasets and improving model interpretability could help bridge the gap between automated ECG diagnosis and real-world clinical deployment.
By advancing the use of deep learning for scanned ECG classification, this work contributes to the broader goal of enhancing accessibility to historical ECG data, supporting telemedicine applications, and enabling automated cardiac diagnostics in resource-constrained settings.